Calculate the osmotic pressure at 37degree celseus. For a solution which has a freezing point of -0.52degree.(kf of water is 1.86)
1
Expert's answer
2020-11-04T14:21:01-0500

ΔTf= Kf x Cm

Cm = ΔTf / Kf = 0.52/1.86 = 0.3 m = 0.3 mol/kg

The osmotic pressure of a solution is the pressure difference needed to stop the flow of solvent across a semipermeable membrane.

Assuming that CM=Cm,

П=C x R x T

П=0.3 x 0.0821 x (37+273)=7.6 atm